OpenStudy (loser66):

x'=Ax where \[A=\left[\begin{matrix}-3&0&2\\1&-1&0\\-2&-1&0\end{matrix}\right]\] I got characteristic equation \[\lambda ^3+4\lambda^2+7\lambda+6=0\\\lambda_1=-2;~\lambda_{2,3}=-1\pm\sqrt{2}i\] continue on comment

OpenStudy (loser66):

for \(\lambda_1 =-2\) I got the eigenvector \(\left(\begin{matrix}2\\-2\\1\end{matrix}\right)\) for \(\lambda_2=-1+\sqrt{2}i\) I got the matrix \[\left[\begin{matrix}-2-\sqrt{2}i&0&2\\1&\sqrt{2}i&0\\-2&-1&1+\sqrt{2}i\end{matrix}\right]\]but don't know how to get the eigenvector from this. Please, help

OpenStudy (loser66):

I step away for awhile, when you stop by, please leave your help. Thanks in advance

OpenStudy (amistre64):

just had to check to make sure your lambdas was good: -2, \(-1\pm i\sqrt2\) are fine

OpenStudy (amistre64):

the next step is to just plug in the lambdas and row reduce to echelon form to define the eugene spaces i believe

OpenStudy (amistre64):

i might have to review that vector step ....

OpenStudy (amistre64):

lol, i used l=2 not -2 on my paper

OpenStudy (amistre64):

1 0 -2 x0 = x3( 2) 0 1 2 x2 = x3(-2) so (2,-2,1) is an eugene vector .... good 0 0 0 x3 = x3( 1)

OpenStudy (amistre64):

\[\left[\begin{matrix} -3-(-1+i\sqrt2)&0&2\\ 1&-1-(-1+i\sqrt2)&0\\ -2&-1&-(-1+i\sqrt2)\end{matrix}\right]\] \[rref:\left[\begin{matrix} -2-i\sqrt2&0&2\\ 1&-i\sqrt2&0\\ -2&-1&1-i\sqrt2\end{matrix}\right]\]

OpenStudy (amistre64):

\[\left[\begin{matrix} -2-i\sqrt2&0&2\\ 1&-i\sqrt2&0\\ -2&-1&1-i\sqrt2\end{matrix}\right]\] \[\left[\begin{matrix} 1&-i\sqrt2&0\\ -2&-1&1-i\sqrt2\\ -2-i\sqrt2&0&2\\ \end{matrix}\right]\] \[\left[\begin{matrix} 1&-i\sqrt2&0\\ 0&-1-2i\sqrt2&1-i\sqrt2\\ 0&2-2i\sqrt2&2\\ \end{matrix}\right]\] \[\left[\begin{matrix} 1&-i\sqrt2&0\\ 0&1&\frac{1-i\sqrt2}{-1-2i\sqrt2}\\ 0&1&\frac{2}{2-2i\sqrt2}\\ \end{matrix}\right]\] \[\left[\begin{matrix} 1&0&i\sqrt2\frac{1-i\sqrt2}{-1-2i\sqrt2}\\ 0&1&\frac{1-i\sqrt2}{-1-2i\sqrt2}\\ 0&0&\frac{2}{2-2i\sqrt2}-\frac{1-i\sqrt2}{-1-2i\sqrt2}\\ \end{matrix}\right]\] etc ...

OpenStudy (amistre64):

might have to wolf it to be sure i didnt get a typing error in there

OpenStudy (amistre64):

it messier, but runniong some conjugates and such is just more time consuming then difficult perse

OpenStudy (loser66):

@amistre64 you mean I have to do the whole stuff to get the rref? WOW!!

OpenStudy (amistre64):

lol, yes ... and you also have to do it correctly to boot

OpenStudy (loser66):

Omg, I thought there is a shortcut. hihihih.....

OpenStudy (amistre64):

shortcut? in linear algebra? ... the only shortcut im aware of is a computer

OpenStudy (loser66):

hahahaaaaa. OK, if I have to, no choice for me. Thanks a lot

OpenStudy (amistre64):

good luck ;)

OpenStudy (loser66):

thank you

OpenStudy (loser66):

let the eigenvector is \[\left(\begin{matrix}a\\b\\c\end{matrix}\right)\] from the matrix above , I have\[\left[\begin{matrix} 1&-i\sqrt2&0\\ -2&-1&1-i\sqrt2\\ -2-i\sqrt2&0&2\\ \end{matrix}\right]\rightarrow \left[\begin{matrix} 1&-i\sqrt2&0\\ 0&-1-2\sqrt{2}i&1+\sqrt{2}i\\ 1&0&\frac{1}{6}(2-\sqrt{2}i)\\ \end{matrix}\right]\] To get it, I do : 2*row1 + row 2 to get row 2 and 1/6 *(2-sqrt(2) i) *row3 to get row 3 \[\text{My question is: can I step up by doing this}\] \[a = i\sqrt{2}b\\a=\frac{1}{6}(2-\sqrt{2}i)c \]and my \[\left(\begin{matrix}a\\b\\c\end{matrix}\right)=\left(\begin{matrix}1\\\frac{-i\sqrt{2}}{2}\\2+\sqrt{2}i\end{matrix}\right)\] @amistre64